GlueX Start Counter Meeting, June 12, 2014

From GlueXWiki
Jump to: navigation, search

GlueX Start Counter Meeting
Thursday, June 12, 2014
10:00 am EDT


  1. Announcements
  2. Review minutes from previous meeting
  3. Detector Construction Status (Eric)
  4. Support Structure Construction Status (Tim/Chuck)
  5. Schedule Review (all)
  6. Software Status (all)
    1. Bias Voltage Control: planned for August

Meeting Information

JLab Meeting Room: CEBAF Center F326


The BlueJeans Meeting ID is 556286544 .

  • Web Browser
  • Room System (e. g. Polycom H.323 device)
    1. or
    2. Enter Meeting ID or use the pairing code shown on the room system
  • Telephone
    1. Dial-in: +1 408 740 7256 or +1 888 240 2560 (US or Canada only) (other phone numbers)
    2. Enter Meeting ID


Talks can be deposited in the directory /group/halld/www/halldweb/html/talks/2014-2Q on the JLab CUE. This directory is accessible from the web at .


  • FIU: Werner Boeglin, Eric Pooser
  • JLab: Eugene Chudakov, Chuck Hutton, Mark Ito (chair), Bruce Lenzer, Tim Whitlatch, Beni Zihlmann

Review of Minutes from Previous Meeting

We reviewed the minutes from the May 29th meeting.

  • Eric reported that Nick Sandoval has all the cables and the MPOD crated and ready to ship. Nick is still doing the individual MPPC tests with the laser, masking non-participating MPPCs.

Detector Construction Status

Eric reviewed a problem with the polishing on the edges of the scintillator.

  • Looking at the edges with wrapping. The protective film might have messed up view. It was removed on 17 of the counters.
  • He saw buffing scratches in the same location on all counters
  • Some chips on the edges, although there were also some of these on the prototypes.
  • Polishing was not done on tip in some cases.
  • He shined a laser into the counters and looked with UV light for bubbles.
    • no crazing
    • no bubbles
    • all defects on edges
  • The light tack adhesive was not removed at FIU on those beyond the first 17.
  • The scintillator was returned to McNeal for polishing of the edges.
  • Bruce contacted the company to discuss our concerns with their process.
  • The counters arrived back yesterday, and visual inspection of the first few show good edge surface quality.
  • Tim mentioned that we should add a note about the polishing to the mechanical drawings.
  • Initial tests of the scintillators at the SiPM end show time resolution 75 to 100 ps worse than seen before. The suspicion is that in the studying light transmission across the air gap with various configurations the surface of the SiPM go damaged. Eric has shipped the SiPM board back to Nick Sandoval. Nick will have a look and send back an up-to-date version of the board as a replacement. Testing is thus on hold for a few days.
  • The assembly jig is all ready to go apart from the support structure.

[Added in press] Further inspection of the scintillators shows area of scratches in two locations on many of the scintillators that were not there before they were returned for edge polishing. In addition, Eric was told that McNeal had done cleaning of the wide surfaces with a commercial glass cleaner in order to remove buffing compound. We will meet tomorrow (Friday) to discuss implications.

Support Structure Status

Tim and Chuck reported that the support structure is complete. It is in the Survey and Alignment Lab to check dimensions and configuration. The shipping case is ready; care was taken to minimize the possibility of damage en route. It should ship to FIU early next week.

Readiness Review

We discussed what is expected for the review. Bruce will draw up a charge to the committee and send that to Werner. We decided on July 8 as a target date.

Software Status

  • Eric has been working on developing the method for time-walk corrections for the past few weeks. It may be that we will need different constants for the different regions of the counter: straight, bend, and nose. He is getting time resolution using a leading-edge discriminator within 10 to 15 ns of that obtained with a constant-fraction discriminator.
  • Werner thought that the geometry currently in HDDS is very close to correct. He has code to generate HDDS from understood parameters. He will look into this area.
  • The start counter is currently being used in the reconstruction.
  • Mark spoke to Hovanes Egiyan about plans for bias voltage control. A GUI will be available in August.

Light Tightening the Start Counter

We discussed several ideas.

  • black tape or Tedlar against the edges of the electronic boards
  • plastic ring on the outside of the boards
  • silicon caulk to fill vulnerable spaces

This is likely not an insurmountable problem, but requires more thinking and discussion.